OpenStudy (anonymous):

Skid mark analysis is one way to estimate the speed a car was traveling prior to an accident. The speed, s, that the vehicle was traveling in mph can be approximated by using the formula s=√30df, where d represents the length of the skid marks in ft and f represents the drag factor of the road. A vehicle that was in an accident on an asphalt road was traveling at 50 mph when it started skidding. If the drag factor for asphalt is 0.75, find the length of the skid marks made by the car. ?? feet long?

OpenStudy (anonymous):

50=sqrt(30*x*.75) 2500=22.5x x=111

OpenStudy (dumbcow):

plug in what you know s = 50 f = 0.75 solve for d 50 = sqrt(30)*d*0.75

OpenStudy (anonymous):

if the sqrt engulfs everything beside it.

OpenStudy (dumbcow):

oh yeah you are right no way it can 12 ft